22. Given: C = 400 + 6(Y) a. Calculate Cify - $100....if Y-S1200?...if Y-$3000? b. Calculate the MPC. Calculate the MPS.
Answer #1

Answer

a)

Y=100

C=400+0.6*Y=400+0.6*100=$460

the consumption is $460

Y=1200

C=400+0.6*Y=400+0.6*1200=$1120

the consumption is $1120

Y=3000

C=400+0.6*Y=400+0.6*3000=$2200

the consumption is $2200

b)

MPC=marginal propensity to consume =coefficient of Y in the consuption funtion =0.6

MPS=1-MPC=1-0.6=0.4